Suppose a project requires an initial investment of $2000 and it is expected to generate a



cash flow of $100 for 3 years plus $12500 in the third year. The target rate of return of



the project is 10% per annum. Calculate the net present value of the project.


1
Expert's answer
2022-01-10T05:13:27-0500

We shall find the net present value for the project after 3 years

we know the formula for net present value is

(ALL CALCULATIONS IN $)

"\sum_{t=0}^{n}\frac{cash flow}{(1+i)^{t}}-initial" "investment"

we let "R_{t}=cash\hspace{0.1cm}flow" at point "t"

"\sum_{t=0}^{n}\frac{cash flow}{(1+i)^{t}}"


"\sum_{t=0}^{n}\frac{R_{t}}{(1+i)^{t}}"

where "n=3"

"=\frac{100}{(1+0.10)^{0}}+\frac{100}{(1+0.10)^{1}}+\frac{100}{(1+0.10)^{2}}+\frac{12500}{(1+0.10)^{3}}"


"=\frac{100}{(1.10)^{0}}+\frac{100}{(1.10)^{1}}+\frac{100}{(1.10)^{2}}+\frac{12500}{(1.10)^{3}}"


"=90.91+82.64+9391.44"

"=9564.99"


NOW,

The Net Present Value is thereby given by

"NPV=9564.99-initial\hspace{0.1cm}investment"

"=9564.99-2000\\=7564.99"

Therefore, the Net Present Value at the end of the third year is $"7564.99"